Hawaii-2017-02 experiment with new MOBYrefresh blue Resonon - BS03cfg001

BS03cfg001 (cfg001 is Resonon data) which may go on the MOBY263 deployment. According to an email on day 5 from Mike "The BS03 is not going to deploy with MOBY263, but rather it is going to get configured with the BS housing from M261 and sent to Miami for long-term testing purposes". In telcon Mike said he was wrong the BS03 will be going in MOBY263 and it will not be sent to Miami but mocked up in Hawaii for Art to do testing of data aquisition. See Hawaii-2017-05 for pre deployment data with full optical setup

Problems seen (and solved) and good things:
  1. On day2 the higher integration time backgrounds have a hump on most of the tracks between pixel 700 - 950 and there is an odd curved feature on track 1. The one dark file does not shows the hump but still has the track 1 curved feature.

Log Sheets

Photos

2nd Order Polynomial fit parametes for the 14 wavelengths - day 3
P(01,:) = [1.080525e-05     0.3403261      337.1167]  
P(02,:) = [1.087847e-05     0.3401773       337.179]  
P(03,:) = [1.07523e-05     0.3402973      337.1438]  
P(04,:) = [1.091682e-05     0.3400416      337.2388]  
P(05,:) = [1.082988e-05     0.3401418      337.2083]  
P(06,:) = [1.094241e-05     0.3399981      337.2884]  
P(07,:) = [1.096924e-05     0.3399546       337.327]  
P(08,:) = [1.083013e-05     0.3401361      337.2339]  
P(09,:) = [1.09482e-05     0.3400136      337.2893]  
P(10,:) = [1.084522e-05     0.3402042      337.2109]  
P(11,:) = [1.096868e-05     0.3400797      337.2559]  
P(12,:) = [1.089064e-05     0.3401832      337.1952]  
P(13,:) = [1.113427e-05     0.3399417      337.2704]  
P(14,:) = [1.098753e-05     0.3401516       337.193]  

 

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 23 Mar Raw Data , Table
DAY01 - bare fiber Oriel 6000 - check track definitions
Mar 23, 2017
1.02
Day 02 - 25 Mar Raw Data , Table
DAY02 - bare fiber Oriel 6000 - looking at saturation behaviour
Mar 27, 2017
1.03
Day 03 - 29 Mar Raw Data , Table
DAY03 - bare fiber - wave cal with Ar, HgA, Ne, Kr, Xe lamps
Mar 28, 2017
1.04
Day 04 - 31bMar Raw Data , Table
DAY04 - bare fiber - laser data
Mar 31, 2017
1.05
Day 05 - 4 Apr Raw Data , Table
DAY05- bare fiber - more laser data and repeats
Apr 4, 2017
 
Day01 - bare fiber Oriel 6000 - check track definitions
2.02 Checking tracks Looking at track definitions Mar 23, 2017
Day02 - bare fiber Oriel 6000 - looking at saturation behaviour
3.01 Sat Check Looking at saturation behavious Mar 27, 2017
3.02 Backgrounds/Darks Looking at backgrounds and darks (features seen in backgrounds??) Mar 27, 2017
3.03 Checking saddle shapes Mike says "I'm wondering how consistent the saddle shape stays as one fills up the ADU counter?" Mar 27, 2017
Day03 - bare fiber - wave cal with Ar, HgA, Ne, Kr, Xe lamps
4.01 Rough wavecal Rough wavecal using 5 lamps (Ar, HgA, Ne, Kr, Xe lamps) Mar 28, 2017
DAY04 - bare fiber - laser data
5.01 Laser data First set of laser data on BS03 and first since the move to Pier 35 Mar 28, 2017
DAY05- bare fiber - more laser data and repeats
6.01 Laser data Second set of laser data on BS03 Apr 4, 2017
6.02 Laser data - details Looking a laser repeats and Mikes dark experiment (all looks good) Apr 4, 2017
       
Multi day graphs
100.01 Saddle shapes Looking at saddle shape with different sources - normed to max Apr 6, 2017
100.02 Saddle shapes V2 Looking at saddle shape with different sources - normed to mean Apr 7, 2017
       

 

DAY 1 EMAIL from Mike

Hi Steph,
Ready or not, here we go!
I have HI-2017-02 day01 data for thee for at /ftp1/Mike/HI-2017-02. This is the BS03cfg001 which may go on the MOBY263 deployment. There is 1 set of sig/drk/bac .fits files to check the track alignment. There were no Aux data collected, but I took some photos, and there are a pair of data log sheets.  The *log01a*.jpg log sheet has a list of the misc. parameters from the SOLIS software's "about" page.
Luck, MF

DAY 2 EMAIL from Mike

Hi Stephs,
I got a quick set of BS03 all-tracks bare-fibers Oriel-incandescent data to look at saturation behaviour: 13x sig,bac scans from .05 to 14 sec, plus 1x dark scan @ 14 sec.  There is 1x log sheet, no photos, no aux data.

...I was looking at the day01 data you processed and how the signal wings seem to not come all the way down to background level between tracks...

I did setup the aux data acq and took some test data which I'll have a look at (and try to remember how to process it!) before I start including that.
Happy Weekend! MF

DAY 3 EMAIL from Mike

Hi StEph,
I collected some wave cal data today via BS03cfg001. There were 3x pairs of sig,bac for each of 5x lamps = Ar, HgA, Ne, Kr, Xe.  It looks like you might have last run through this data set in Apr-2017, for HI-2017-02, BS01cfg003. If you have any confusion r.e. the wave lamps then I can dig out my MOS notes, because they are always confusing to I!

The HgA lamp can be trouble because there are VERY intense near-UV Hg lines at 253.6521 & 296.7283 nm, which MOS sees as second order @ *2 = 507.3042 & 593.4566 nm - I'm very curious if this is also so for the BS?  There are also 3x more near-UV lines in Hg at 312.5674, 313.155, 314.1844 that may be seen doubled at 625.1348, 626.31, 628.3688 nm, but they are not as bright as the 253 & 296 lines.

I did run the BS aux data collection today = 1x aux*.dat file, plus there was 1x log sheet, and I took 3x photos - the photos show the Oriel pen lamp setup with the 12 in Spectralon sphere, where we have an inside-the-sphere mount for the pen lamps, but I had to mount the HgA & Ne lamps outside the sphere
because they were too dern bright.
Luck, MF

DAY 4 EMAIL from Mike

Hello again,
I started collecting some laser data with the BS03cfg001 = HI-2017-02, day04, GMT 31-Mar-2017, with 2x log sheets, 1x aux file, and what should have been 18x sig,bac .fits pairs. I seemed to have has some trouble saving files - both the sig & bac were not saved for *08 & *09.fits, and the bac was not saved for *17.fits. I apologize for the headaches that will cause!

These were the first laser data since the Snug/Pier35 move so it will be interesting to look for any sign of odd lasering, but for the most part the Ekspla seemed to work OK (once I remembered how to turn everything on). This was a VERY coarse (sub)set of wavelengths.

Happy Cesar Chavez Day!  MF

DAY 5 EMAIL from Mike
Hi Steph,
Well, there was a change of plans today, that, or perhaps I misunderstood what I thought was the ole plan.

The BS03 is not going to deploy with MOBY263, but rather it is going to get configured with the BS housing from M261 and sent to Miami for long-term testing purposes.

Today the BS02 was removed from the M261 housing so it's camera CCD-16345 could be sent back to Andor for repairs.

So, there is not a great need to continue characterizing the BS03 now, but, since we were at it and all was set up, I decided to get another set of laser scans.  These day05 are some 5x sig,bac multiples of scans from day04, plus some fill-in lines at the second-order kine, plus some fill-ins @ 25 nm spacing across the array, plus some off-ccd lines on the red-side.  There was 1x aux file and 3x log sheets.

I also collected 2x dark scans.  The first was darkened by disconnecting the lasers' FO from the sphere.  The second was darkened by re-connecting the FO @ the sphere and setting the pump-laser to off. The difference between the two darks (and the first 5x sig,bac @ 340 nm) would be that drk #2 had the possibility that ambient light between the laser & the laser's FO coupler could possibly get into the sphere via the FO. I think/hope dark #2 is "dark".

There was also a "slight" laser config change between scan #1 & #2. For scan #1, and all of day04 data, the yellow "laser on" safety light on the end of the laser was not covered.  And, the space between the laser & the laser's FO coupler was open to laptop screen light from your ole MacBook that I use to control the laser.  So, I am really hoping there is no difference between scan #1 & #2...

Thank You, MF